Solve the equation and graph x-4y=8.
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

Solve the equation and graph x-4y=8.
...
x-4y=8
4y=x-8
/4
y=x/4-2
x=0 y=-2
y=0 x=8
...
(0,-2) & (8,0) are the intercepts.
plot the points and draw the line graph
..

...
m.ananth@hotmail.ca

Answer by Alan3354(69443)

A single equation with 2 variables cannot be "solved."
You can solve for x in terms of y, or y in terms of x, but that's all.
The equation is a line, as the other tutor drew. Every point on the line fits the equation.